OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. 2 = x2 + 5x

OpenStudy (imstuck):

Put it into standard form first. Ax^2 + Bx + C = 0.

OpenStudy (anonymous):

x^2+5x=2

OpenStudy (imstuck):

The discriminant is the "inside" part of the quadratic formula, what's under the radical sign.

OpenStudy (imstuck):

Standard form is x^2 + 5x - 2 = 0

OpenStudy (imstuck):

The discriminant formula is b^2 - 4(a)(c). Your a is a 1, your b is a 5, and your c is a -2.

OpenStudy (imstuck):

Fill those values into that formula and see what the value of the discriminant is.

OpenStudy (anonymous):

ok thank you

OpenStudy (imstuck):

\[(5)^{2}-4(1)(-2)=25+8=33\]

OpenStudy (imstuck):

Do you know the rules about the discriminant's values and what it tells you about the graph of the polynomial?

OpenStudy (anonymous):

greather than 0 = rational & real perfect square =0 will be real & rational <0 will be complex or imaginary

OpenStudy (imstuck):

greater than 0, there are 2 real solutions; less than 0 there are no real solutions, only imaginary complex solutions; equals 0, there is 1 real solution.

OpenStudy (imstuck):